# F-95000 Cergy-Pontoise, France

### v1

@@ Abstract
We present an approach to computing consistent answers to analytic queries in data warehouses operating under a star schema and possibly containing missing values and inconsistent data. Our approach is based on earlier work concerning consistent query answering for standard, non-analytic queries in multi-table databases. In that work we presented polynomial algorithms for computing either the exact consistent answer to a standard, non analytic query or bounds of the exact answer, depending on whether the query involves a selection condition or not.
We extend this approach to computing exact consistent answers of analytic queries over star schemas, provided that the selection condition in the query involves no keys and satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute). The main contributions of this paper are: (a) a polynomial algorithm for computing the exact consistent answer to a usual projection-selection-join query over a star schema under the above restrictions on the selection condition, and (b) showing that, under the same restrictions the exact consistent answer to an analytic query over a star schema can be computed in time polynomial in the size of the data warehouse.
Keywords: Star schema $\bullet$ analytic query $\bullet$ inconsistent data $\bullet$ consistent query answering
@@ 1 Introduction
The efficient processing of analytic queries is an important issue in databases and has motivated considerable research work since the last three decades. The main purpose of analytic queries is to extract relevant ‘statistics’ from huge volumes of data, resulting from the integration of heterogeneous databases and stored in what is called a data warehouse inmon96 ; Kimball . For efficiency reasons, the data stored in a data warehouse is generally organized according to a non-normalized schema, called a star schema. A star schema consists of two types of relation schemas (also called  tables): a fact table and a number of dimension tables. In this context, an analytic query can be seen as an SQL Group-by query involving some aggregate function such as $min$, $max$, $count$ or $sum$ operating over attributes in the fact table called measure attributes (or simply measures). Let us see an example to illustrate the concepts of star schema and analytic query.
@@
    • Branch, defined over the attributes $K_{B}$, $B\_Loc$, $B\_Ctry$ standing respectively for the branch identifier, the town in which the branch is located and the country which this town belongs to. The attribute $K_{B}$ is the (primary) key of Branch, meaning that the table Branch must satisfy the functional dependencies $K_{B}\to B\_Loc$ and $K_{B}\to B\_Ctry$.
    • Prod, defined over the attributes $K_{P}$, $P\_Type$, $Price$, where $K_{P}$ is the product identifier, and $P\_Type$ and $Price$ are respectively the type and the price of a product. The attribute $K_{P}$ is the (primary) key of Prod, meaning that the table Prod must satisfy the functional dependencies $K_{P}\to P\_Type$ and $K_{B}\to Price$.
    • Cust, defined over the attributes $K_{C}$, $C\_Name$, $C\_Addr$ standing respectively for customer identifier, name and address of the customer. The attribute $K_{C}$ is the (primary) key of the table Cust, meaning that Cust must satisfy the functional dependencies $K_{C}\to C\_Name$ and $K_{C}\to C\_Addr$.
    • Date, defined over the attributes $K_{D}$, $Month$, $Year$ standing respectively for the date identifier or key, the month and the year of the date. The attribute $K_{D}$ is the (primary) key of Date, meaning that Date must satisfy the functional dependencies $K_{D}\to Month$ and $K_{D}\to Year$.
@@
In this setting a typical analytic query is to display the total quantity of all products sold in all branches during the year $2024$. This query can be expressed in SQL as follows:
@@
Here $J$ denotes the (lossless) join of all dimension tables with the fact table $F$ (although the join can by simplified by involving only $F$, Prod and Date). $\Box$
@@
A more problematic situation is if the selection condition in the query is $Year=2024$ and $C\_Addr=Paris$. This is so because among all transactions regarding customers whose address may be Paris, some concern customers whose address may violate the dependency $K_{C}\to C\_Addr$ in the table Cust. Dealing with such inconsistencies, known as the problem of computing the consistent answer to an analytic query, is not trivial, and as argued in ArenasBCHRS03 ; FuxmanFM05 , techniques used for standard non analytic queries cannot be used for analytic queries.
To cope with inconsistencies and missing values in data warehouses, our approach is based on our earlier work LauS25 dealing with consistent query answering for standard, non analytic queries in multi-table databases. In that work, we presented polynomial algorithms for computing either the exact consistent answer to a standard non analytic query or bounds of the exact answer, depending on whether the query involves or not a selection condition. In the present paper, we show that in the case of a star schema, under the restrictions that the selection condition involves no keys and satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute), the exact consistent answer can be effectively computed. In the following section, we summarize briefly the main results of the approach in LauS25 and then focus on analytic queries over star schemas.
Considering queries, analytic or not, whose selection condition satisfies the two restrictions mentioned above, the main contribution of this paper is showing that: (a) computing the exact consistent answer to a usual projection-selection-join query over a star schema is in time polynomial in the size of the data warehouse (in contrast to LauS25 , where consistent answers to non analytic queries are approximated when involving a selection condition), and (b) the exact consistent answer to an analytic query over a star schema can be computed in time polynomial in the size of the data warehouse (with two exceptions where only approximations are given).
The paper is organized as follows: In Section 2 we recall the main features of our previous work in LauS25 , on which the present approach is based. In Section 3 we first recall the definition of a star schema and argue that the approach in LauS25 applies in this context. In Section 4 we investigate further the concept of repairs in the context of star schemas. Section 5 deals with consistent answers to queries in the case of standard projection-selection queries as well as in the case of analytic queries. In Section 6, we propose algorithms for efficiently computing the consistent answers, or in some few cases, an approximation of the consistent answers to analytic queries. In Section 7 we compare our approach to other approaches from the literature and in Section 8 we summarize the contents of the paper and suggest research directions for future work.
@@ 2 The Approach of LauS25
2.1 The Approach at a Glance
@@
In this context, a tuple $t$ in the current value of $Chase(T)$ is said to be conflicting if the following holds: there is a tuple $t^{\prime}$ in the current value of $Chase(T)$ and a dependency $X\to A$ in $FD$ such that $t$ and $t^{\prime}$ are both defined over $XA$, $t.X=t^{\prime}.X$ and $t.A\neq t^{\prime}.A$. A tuple $t$ is called non-conflicting if $t$ is not a conflicting tuple.
@@
Now, if Chase is successful (i.e., no conflicts are encountered and no more new tuples can be derived) then the database is declared consistent else conflicting. If the database is consistent then processing of queries (whether standard queries or analytic queries) proceeds as usual, else the following question arises: can we still extract useful (i.e. non conflicting) information from the conflicting database? The work in LauS25 gives a positive answer to this question based on an extended version of the Chase algorithm, called the $m\_Chase$ algorithm to be presented formally in the following subsection. The input of this algorithm is no more $T$ but a table $\mathcal{T}$ containing all tuples that can be built up from constants in the active domains of the database. The set $\mathcal{T}$ is called the ‘Universe of discourse’ of the $m\_Chase$ algorithm and the notion of ‘conflicting tuple’ remains the same but now concerns all tuples of $\mathcal{T}$ and not only those of $T$.
@@
It is shown in LauS25 that the tuples of $\mathcal{T}$ can be characterized in two orthogonal ways: a tuple of $\mathcal{T}$ can be either true or false, and it can be either conflicting or non-conflicting. This characterization can be intuitively described as follows. If Chase terminates successfully and $Chase(T)$ denotes the output table then a tuple $t$ of $\mathcal{T}$ is true if it appears in $Chase(T)$ and false otherwise (i.e. $t$ is false if it appears in $\mathcal{T}\setminus Chase(T)$).
@@
However, if the Chase algorithm fails then we don’t know which tuples are true and which are non-conflicting. The $m\_Chase$ algorithm remedies this deficiency by modifying the Chase algorithm as follows: instead of stopping the application of functional dependencies on table $T$ when a conflict is encountered, the application continues (and the true tuples are stored) until no more tuples are found. In doing so, all true tuples and all conflicting tuples are computed - and therefore each tuple of $\mathcal{T}$ can be characterized as true/false and as conflicting/non-conflicting.
@@
It follows from the above definition of conflicting tuple that if $t$ is conflicting then every true super-tuple of $t$ is also conflicting. Therefore the conflicting tuples can be retrieved as true super-tuples of true tuples of the form $xa$ over $XA$ such that: $(a)$ $X\to A$ is a dependency in $FD$ and $(b)$ $a$ and $a^{\prime}$ are in $adom(A)$ such that $a\neq a^{\prime}$ and $xa^{\prime}$ is true (here $adom(A)$ stands for ‘active domain’ of $A$). Then assuming that all true tuples and all conflicting tuples are known, we can define a tuple $t$ of $\mathcal{T}$ to be consistent if $t$ is true and non-conflicting. Note that every sub-tuple of a true tuple is true and that every sub-tuple of a consistent tuple is consistent. Finally, call ‘consistent subset’ of $\mathcal{T}$ any set $S$ of true tuples of $\mathcal{T}$ such that the set of all tuples inferred from $S$ using the functional dependencies contains no conflicting tuples in $\mathcal{S}$. Let us illustrate these concepts using the following example.
@@
Suppose $T=\{ab,bc,ac^{\prime}\}$ and $FD=\{A\to C,B\to C\}$. Then all tuples in $T$ are true and non conflicting (hence consistent), but the application of the functional dependencies on $T$ allows to infer the tuples $abc$ and $abc^{\prime}$, which are conflicting tuples of $\mathcal{T}$ inferred from $T$. In fact it can be seen that $(a)$ the true tuples of $\mathcal{T}$ are $abc$, $abc^{\prime}$ and all their sub-tuples, implying that any other tuple of $\mathcal{T}$ is false in $\mathcal{T}$, and $(b)$ the conflicting tuples in $\mathcal{T}$ are $abc$, $abc^{\prime}$, $ac$, $ac^{\prime}$, $bc$ and $bc^{\prime}$, implying that any other tuple in $\mathcal{T}$ is non conflicting in $\mathcal{T}$. In this example, the consistent tuples of $\mathcal{T}$ are $ab$, $a$, $b$, $c$ and $c^{\prime}$.
@@
In this context, the set $R=\{abc,c^{\prime}\}$ is a consistent subset of $\mathcal{T}$. Indeed, since $\mathcal{R}$ is the set of all tuples built up using constants $a$, $b$, $c$ and $c^{\prime}$, we have $\mathcal{R}=\mathcal{T}$. Moreover the tuples in $R$ are tuples true in $\mathcal{T}$ and no conflicting tuples in $\mathcal{R}$ can be generated from the tuples in $R$. It is important to notice that, although $abc$ is conflicting in $\mathcal{T}$, $abc$ is not conflicting in $\mathcal{R}$ because $ac^{\prime}$, $bc^{\prime}$ and $abc^{\prime}$ are not true in $\mathcal{R}$. $\Box$
@@
Finally, based on the concepts introduced so far, a repair of $T$ in LauS25 is defined to be a maximal and consistent subset of $\mathcal{T}$ containing a maximal and consistent set of tuples which are consistent in $\mathcal{T}$. In our previous example, the subset $R=\{abc,c^{\prime}\}$ of $\mathcal{T}$ is a repair of $T$ because $(a)$ as we have just seen, $R$ is a consistent subset of $\mathcal{T}$, $(b)$ $R$ is maximal because adding to $R$ a tuple true of $\mathcal{T}$ either does not bring any new true tuple in $\mathcal{R}$ (e.g., adding the tuple $ac$) or generates a conflicting tuple in $\mathcal{R}$ (e.g., adding the tuple $ac^{\prime}$), and $(c)$ all consistent tuples of $\mathcal{T}$ are true in $\mathcal{R}$. Note that similar arguments show that the set $S=\{bc,ac^{\prime}\}$ is a maximal and consistent subset of $\mathcal{T}$, however, $S$ is not a repair of $T$, because $ab$ is a consistent tuple of $\mathcal{T}$ which is not true in $\mathcal{S}$. By the way, as we shall see in Section 4.1, our definition of repair is more restrictive than the usual definition ArenasBC99 ; Wijsen19 in which a repair is defined to be a maximal and consistent subset of $\mathcal{T}$. For example, the set $S=\{bc,ac^{\prime}\}$ is a repair of $T$ following ArenasBC99 ; Wijsen19 , but it is not a repair of $T$ following our approach.

### v2

@@ Abstract
We present an approach to computing consistent answers to queries possibly involving an aggregation operator in databases operating under a star schema and possibly containing missing values and inconsistent data. Our approach is based on earlier work concerning consistent query answering for standard queries (with no aggregate operator) in multi-table databases. In that work, we presented polynomial algorithms for computing either the exact consistent answer to a query or bounds of the exact answer, depending on whether the query involves a selection condition or not.
In the present work, we consider databases operating under a star schema. Calling data warehouses such databases, we extend our previous work to queries involving aggregate operators, called analytic queries. In this context, we propose specific algorithms for computing exact consistent answers to queries, whether analytic or not, provided that the selection condition in the query satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute). We show that the overall time complexity of these specific algorithms is in $O(W\cdot\log(W))$, where $W$ is the size of the data warehouse. Moreover, the case of analytic queries involving a having clause associated with a group-by clause is discussed in the context of our approach.
Keywords: Star schema, analytic query, inconsistent data, consistent query answering
@@ 1 Introduction
The efficient processing of queries involving an aggregation operator, hereafter referred to as analytic queries, is an important issue in databases and has motivated considerable research work since the last three decades. The main purpose of analytic queries is to extract relevant ‘statistics’ from huge volumes of data, resulting from the integration of heterogeneous databases and stored in what is called a data warehouse inmon96 ; Kimball . For efficiency reasons, the data stored in a data warehouse is generally organized according to a non-normalized schema, called a star schema. A star schema consists of two types of relation schemas (also called tables): a fact table and a number of dimension tables. In this context, an analytic query can be seen as an SQL Group-by query involving some aggregate operator such as minmin, maxmax, countcount or sumsum operating over attributes in the fact table called measure attributes (or simply measures). Let us see an example to illustrate the concepts of star schema and analytic query.
@@
    • Branch, defined over the attributes $K_{B}$, B_TownB\_Town, B_CtryB\_Ctry standing respectively for the branch identifier, the town in which the branch is located and the country which this town belongs to. The attribute $K_{B}$ is the (primary) key of Branch, meaning that the table Branch must satisfy the functional dependencies $K_{B}\to B\_Town$ and $K_{B}\to B\_Ctry$.
    • Prod, defined over the attributes $K_{P}$, P_TypeP\_Type, PricePrice, where $K_{P}$ is the product identifier, and P_TypeP\_Type and $Price$ are respectively the type and the price of a product. The attribute $K_{P}$ is the (primary) key of Prod, meaning that the table Prod must satisfy the functional dependencies $K_{P}\to P\_Type$ and $K_{B}\to Price$.
    • Cust, defined over the attributes $K_{C}$, C_NameC\_Name, C_AddrC\_Addr standing respectively for customer identifier, name and address of the customer. The attribute $K_{C}$ is the (primary) key of the table Cust, meaning that Cust must satisfy the functional dependencies $K_{C}\to C\_Name$ and $K_{C}\to C\_Addr$.
    • Date, defined over the attributes $K_{D}$, $Month$, YearYear standing respectively for the date identifier or key, the month and the year of the date. The attribute $K_{D}$ is the (primary) key of Date, meaning that Date must satisfy the functional dependencies $K_{D}\to Month$ and $K_{D}\to Year$.
@@
In this setting a typical analytic query is to display the total quantity of each product sold during the year $2024$. This query can be expressed in SQL as follows:
@@
Here $J$ denotes the (lossless) join of all dimension tables with the fact table $F$ (although the join can be simplified by involving only $F$, Prod and Date). $\Box$
@@
A more problematic situation is if the selection condition in the query is $Year=2024$ and C_Addr=ParisC\_Addr=Paris. This is so because among all transactions regarding customers whose address may be Paris, some concern customers whose address may violate the dependency $K_{C}\to C\_Addr$ in the table Cust. Dealing with such inconsistencies, known as the problem of computing the consistent answer to an analytic query, is not trivial, and as argued in ArenasBCHRS03 ; FuxmanFM05 , techniques used for standard non analytic queries cannot be used for analytic queries.
To cope with inconsistencies and missing values in data warehouses, our approach is based on our earlier work LauS25 dealing with consistent query answering for standard, non analytic queries in multi-table databases. In that work, we presented polynomial algorithms for computing either the exact consistent answer to a standard non analytic query or bounds of the exact answer, depending on whether the query involves a selection condition or not. In the present paper, we show how to compute effectively the exact consistent answer in the case of a star schema, under the restriction that the selection condition satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute).
The main contributions of this paper can be summarized as follows: We propose specific algorithms for computing exact consistent answers to queries, whether analytic or not, over a star schema, provided that the selection condition in the query satisfies the property of independency. We also show that the overall time complexity of our specific algorithms is in $O(W\cdot\log(W))$, where $W$ is the size of the data warehouse. We discuss the case of analytic queries involving a having clause associated with a group-by clause in the context of our approach.
The paper is organized as follows: In Section 2 we recall the main features of our previous work in LauS25 , on which the present approach is based. In Section 3 we first recall the definition of a star schema and argue that the approach in LauS25 applies in this context. In Section 4 we investigate the concept of repairs in the context of star schemas. Section 5 deals with consistent answers to queries in the case of standard projection-selection queries as well as in the case of analytic queries. In Section 6, we propose algorithms for efficiently computing the consistent answers to analytic queries, or in some few cases an approximation to the consistent answer. In Section 7 we compare our approach to other approaches from the literature and in Section 8 we summarize the contents of the paper and suggest research directions for future work.
@@ 2 The Approach of LauS25
2.1 Basic Definitions and Notation
As done in JIIS ; LauS25 , we rely on basic notions from relational databases as introduced in the literature Ullman . As usual in this context, we consider a finite set of attributes $U=\{A_{1},\ldots,A_{n}\}$, called the universe, and we assume that each attribute $A_{i}$ is associated with a domain of values denoted $dom(A_{i})$. A nonempty subset $R=\{A_{i_{1}},\ldots,A_{i_{p}}\}$ of $U$ is called a (relation) schema and simply denoted by $A_{i_{1}}\ldots A_{i_{p}}$ (i.e., by the concatenation of its attributes). Given such a schema $R=A_{i_{1}}\ldots A_{i_{p}}$, a tuple $t$ over $R$ is an element of the cross product $dom(A_{i_{1}})\times\ldots\times dom(A_{i_{p}})$, and $t$ is denoted by the concatenation of its values, that is $t=a_{i_{1}}\ldots a_{i_{p}}$ where $a_{i_{j}}$ is in $dom(A_{i_{p}})$, for every $j=1,\ldots,p$. In this case, $R$ is called the schema of $t$ and is also denoted by $sch(t)$. Moreover, if $S$ is a subset of $R$, the restriction of $t$ over $S$, denoted by $t.S$ is the tuple over $S$ whose values over attributes in $S$ are those of $t$. More formally, given tuple $t$ over $R$, $S\subseteq R$, and tuple $q$ over $S$, the fact that $q=t.S$ means that $sch(q)=S$ and that for every $A\in S$, $q.A=t.A$.
In this setting a relation, also called a table, over schema $R$ is a finite set of tuples over $R$, and a multi-table database over universe $U$ is a set of such relations. Additionally, relations in a database are associated with constraints, whose goal is to restrict their content so as to model given properties. Functional dependencies, that we consider in our work, are recognized as the most popular constraints.
Given a relation $r$ over schema $R$, a functional dependency over $R$ (or simply ‘dependency over $R$’) is an expression of the form $X\to Y$ where $X$ and $Y$ are subsets of $R$. A relation $r$ over $R$ is said to satisfy the dependency $X\to Y$ if for all tuples $t$ and $t^{\prime}$ in $r$ the following holds: if $t.X=t^{\prime}.X$ then $t.Y=t^{\prime}.Y$; and $r$ is said to satisfy a set $FD$ of dependencies over $R$ if $r$ satisfies every dependency of $FD$. A basic result from functional dependency theory Armstrong says that: $r$ satisfies a set $FD$ of dependencies over $R$ if and only if $r$ satisfies the set $FD^{\prime}=\{X\to A~|~X\to Y\in FD,A\in Y\}$. For example, if $r$ is a relation over $R=ABCDE$ then $r$ satisfies $FD=\{AB\to CD,D\to EC\}$ if and only if $r$ satisfies $FD^{\prime}=\{AB\to C,AB\to D,D\to E,D\to C\}$. This explains why in our work we restrict our attention to dependencies having a right hand-side consisting of just one attribute. Next, we address the issue of checking whether a given multi-table database satisfies a given set of dependencies.
2.2 From $Chase$ to m_Chasem\_Chase
@@
When running the Chase algorithm, a tuple $t$ in the current value of $Chase(T)$ is said to be conflicting if the following holds: there is a tuple $t^{\prime}$ in the current value of $Chase(T)$ and a dependency $X\to A$ in $FD$ such that $t$ and $t^{\prime}$ are both defined over $XA$, $t.X=t^{\prime}.X$ and $t.A\neq t^{\prime}.A$. A tuple $t$ is called non-conflicting if $t$ is not a conflicting tuple.
@@
Now, if Chase is successful (i.e., no conflicts are encountered and no more new tuples can be derived) then the database is declared consistent else conflicting. If the database is consistent then processing of queries (whether standard queries or analytic queries) proceeds as usual, else the following question arises: can we still extract useful (i.e., non conflicting) information from the conflicting database? The work in LauS25 gives a positive answer to this question based on an extended version of the Chase algorithm, called the m_Chasem\_Chase algorithm (to be presented formally in the following subsection). As will shall see, the m_Chasem\_Chase algorithm operates over the set of all tuples that can be built up from constants from the active domains of the database, rather than over the tuples in $T$. We recall that the active domain of an attribute $A$ in a multi-table database is the set of all values of the domain of $A$ that are present in the database. Note that every tuple appearing in the database is built up from constants in the active domains of database attributes but there are tuples that are built-up from constants in the active domains that are not present in the database. For example, if a database relation has only two tuples, $ab$ and $a^{\prime}b^{\prime}$, then the tuples $ab^{\prime}$ and $a^{\prime}b$ are built up from constants in the active domains (here $\{a,a^{\prime}\}$ and $\{b,b^{\prime}\}$) but they don’t belong to the relation.
In order to deal with this important difference between Chase and m_Chasem\_Chase, we introduce the following notation: given a set of tuples $S$, we denote by $\mathcal{S}$ the set of all tuples that can be built up from constants occurring in $S$. In our previous example, where $S=\{ab^{\prime},a^{\prime}b\}$, we have $\mathcal{S}=\{ab,ab^{\prime},a^{\prime}b,a^{\prime}b^{\prime},a,a^{\prime},b,b^{\prime}\}$. It is important to note that when considering the table $T$, the set $\mathcal{T}$ is precisely the set of all tuples that can be built up from constants in the active domains of attributes in $U$. Hereafter, we shall refer to $\mathcal{T}$ as the universe of discourse of the m_Chasem\_Chase. To summarize, the $Chase$ algorithm operates over the set $T$ of all tuples in the database, whereas the m_Chasem\_Chase algorithm operates over the set $\mathcal{T}$. Moreover, in the m_Chasem\_Chase algorithm the notion of ‘conflicting tuple’ remains the same as for the Chase algorithm, but now this notion concerns all tuples of $\mathcal{T}$, not just those of $T$. More precisely, it is shown in LauS25 that the tuples of $\mathcal{T}$ can be characterized in two orthogonal ways: a tuple of $\mathcal{T}$ can be either true or false, and it can be either conflicting or non-conflicting. This characterization can be intuitively described as follows:
@@
  • If Chase algorithm terminates successfully and $Chase(T)$ denotes the output table, then a tuple $t$ of $\mathcal{T}$ is true if it appears in $Chase(T)$ and false otherwise (i.e., $t$ is false if it appears in $\mathcal{T}\setminus Chase(T)$). In this case no tuple in $\mathcal{T}$ is conflicting, entailing that all tuples in $\mathcal{T}$ are non-conflicting.
@@
  • However, if the Chase algorithm fails then we don’t know which tuples are true and which are non-conflicting. The m_Chasem\_Chase algorithm remedies this deficiency by modifying the Chase algorithm as follows: instead of stopping the application of functional dependencies on table $T$ when a conflict is encountered, the application continues (and the true tuples are stored) until no more tuples are found. In doing so, all true tuples and all conflicting tuples are computed - and therefore each tuple of $\mathcal{T}$ can be characterized as true/false and as conflicting/non-conflicting.
@@
It follows from the above definition of conflicting tuple that if $t$ is conflicting then every true super-tuple of $t$ is also conflicting. Therefore the conflicting tuples can be retrieved as true super-tuples of true tuples of the form $xa$ over $XA$ such that: $(a)$ $X\to A$ is a dependency in $FD$ and $(b)$ $a$ and $a^{\prime}$ are in $adom(A)$ such that $a\neq a^{\prime}$ and $xa^{\prime}$ is true (here $adom(A)$ stands for ‘active domain’ of $A$). Then assuming that all true tuples and all conflicting tuples are known, we can define a tuple $t$ of $\mathcal{T}$ to be consistent if $t$ is true and non-conflicting. Note that every sub-tuple of a true tuple is true and that every sub-tuple of a consistent tuple is consistent. Then, a set $S$ of true tuples of $\mathcal{T}$ is said to be a consistent subset of $\mathcal{T}$ if the set of all tuples inferred from $S$ using the functional dependencies contains no conflicting tuples in $\mathcal{S}$. Let us illustrate these concepts using the following example.
@@
Consider a universe of three attributes $A$, $B$ and $C$ over which the functional dependencies $FD=\{A\to C,B\to C\}$ are assumed to hold, along with a database consisting of three relations, namely $r_{1}=\{ab\}$, $r_{2}=\{bc\}$ and $r_{3}=\{ac^{\prime}\}$, respectively defined over schemas $AB$, BCBC and $AC$. Since no dependency can be enforced on $r_{1}$, and since $r_{2}$ and $r_{3}$ respectively satisfy $B\to C$ and $A\to C$, the database satisfies the dependencies in $FD$.
@@
However, when ‘chasing’ this database, its tuples are first collected in a single table $T$ defined over the full universe $ABC$ and containing the three tuples $\{ab,bc,ac^{\prime}\}$, in which missing values occur (for example, $ab$ is not defined over attribute $C$). At this stage, all tuples in $T$ are true (since they are in the database) and they are non conflicting (since the dependencies in $FD$ are satisfied). However, the application of these functional dependencies on $T$ allows to infer the tuples abcabc and $abc^{\prime}$, which are conflicting tuples of $\mathcal{T}$ inferred from $T$. In fact it can be seen that $(a)$ the true tuples of $\mathcal{T}$ are abcabc, $abc^{\prime}$ and all their sub-tuples, implying that any other tuple of $\mathcal{T}$ is false in $\mathcal{T}$, and $(b)$ the conflicting tuples in $\mathcal{T}$ are abcabc, $abc^{\prime}$, $ac$, $ac^{\prime}$, $bc$ and $bc^{\prime}$, implying that any other tuple in $\mathcal{T}$ is non conflicting in $\mathcal{T}$. In this example, the consistent tuples of $\mathcal{T}$ (i.e., the tuples that are true and non conflicting in $\mathcal{T}$) are $ab$, $a$, $b$, $c$ and $c^{\prime}$.
@@
In this context, for $R=\{abc,c^{\prime}\}$, we denote by $\mathcal{R}$ the set of all tuples that can be built up using the constants occurring in $R$, namely $a$, $b$, $c$ and $c^{\prime}$. Thus, we have $\mathcal{R}=\mathcal{T}$ and it can be seen that applying the $Chase$ algorithm to the tuples in $R$ does not change $R$, and does not fail. Therefore, no tuples of $\mathcal{R}$ are conflicting, and since $R$ is a set of tuples true in $\mathcal{R}$, $R$ is said to be a consistent subset of $\mathcal{T}$. We notice that although abcabc is conflicting in $\mathcal{T}$, abcabc is not conflicting in $\mathcal{R}$ because $ac^{\prime}$, $bc^{\prime}$ and $abc^{\prime}$ are not true in $\mathcal{R}$. $\Box$
@@
Based on the concepts introduced so far, a repair of $T$ in LauS25 is defined to be a maximal and consistent subset of $\mathcal{T}$ containing a maximal and consistent set of tuples which are consistent in $\mathcal{T}$. In our previous example, the subset $R=\{abc,c^{\prime}\}$ of $\mathcal{T}$ is a repair of $T$ because $(a)$ as we have just seen, $R$ is a consistent subset of $\mathcal{T}$, $(b)$ $R$ is maximal because adding to $R$ a tuple true of $\mathcal{T}$ either does not bring any new true tuple in $\mathcal{R}$ (e.g., adding the tuple $ac$) or generates a conflicting tuple in $\mathcal{R}$ (e.g., adding the tuple $ac^{\prime}$), and $(c)$ all consistent tuples of $\mathcal{T}$ are true in $\mathcal{R}$. Note that similar arguments show that the set $S=\{bc,ac^{\prime}\}$ is a maximal and consistent subset of $\mathcal{T}$, however, $S$ is not a repair of $T$, because $ab$ is a consistent tuple of $\mathcal{T}$ which is not true in $\mathcal{S}$.
By the way, as we shall see in Section 4.1, our definition of repair is more restrictive than the usual definition ArenasBC99 ; Wijsen19 in which a repair is defined to be a maximal and consistent subset of tuples true in $\mathcal{T}$. In the context of the previous example, since the set $S=\{bc,ac^{\prime}\}$ is a maximal and consistent subset of tuples true in $\mathcal{T}$, $S$ is seen as a repair of $T$ following ArenasBC99 ; Wijsen19 , whereas $S$ is not a repair of $T$ following our approach, because $ab$ is consistent in $\mathcal{T}$ but not true in $\mathcal{S}$.